OpenStudy (anonymous):

For problem 1, I understand that I have to find g = gf*gi^-1 first in order to find xi. How do I put gf and gi into homogenous form for this calculation?

OpenStudy (anonymous):

gf and gi are given in homogeneous form in the problem. Remember homogeneous representation is just the matrix representation of frames. The opposite of that would be the vector representation which would be \[\left(\begin{matrix}x \\ y \\ z\\ \alpha_x\\ \alpha_y\\a_z\end{matrix}\right)\]

OpenStudy (anonymous):

So if g = [R d; 0 1] for homogeneous form, what is my R and D from the given matrices?

OpenStudy (anonymous):

R is a 3x3 matrix and d is a 3x1 vector.

OpenStudy (anonymous):

o wow, i didn't realize it was that simple

OpenStudy (anonymous):

Yep pretty nice

OpenStudy (anonymous):

xi = [w; v]. w = atan2(R21, R22). v = wJ(1-R)^-1*d Is this correct?

OpenStudy (anonymous):

I'm not sure what problem exactly you're talking about but w does not equal atan2(R21,R22). That would be a way to extract the angle from a rotation matrix in the 2D case.

OpenStudy (anonymous):

Whereas w is an angular velocity.

OpenStudy (anonymous):

v = |w|^2*[(1-R)w+ww^T*tau]^-1*d How do I find w?

OpenStudy (anonymous):

Which problem are you working on?

OpenStudy (anonymous):

I'm still trying to calculate xi for problem 1.

OpenStudy (anonymous):

You can extract w from w^